OpenStudy (anonymous):

3x + 4 = 10, then 3x = 6. This is an illustration of which property? Commutative Subtraction property of equality Multiplication property of equality Substitution

OpenStudy (anonymous):

so you know how it went from \[3x+4=10\] so \[3x=6\]?

OpenStudy (anonymous):

substitution?

OpenStudy (anonymous):

actually the step was to subtract 4 from both sides

OpenStudy (anonymous):

\[3x+4=10\] \[3x+4-4=10-4\] \[3x=6\]so it is the "subtraction property"